OpenStudy (anonymous):

verify cscxsecx=2csc2x please help me!!!!!!!

OpenStudy (arthur_ser):

cscxsecx=2csc2x (1/sinx)(1/cosx)=2(1/sin2x) sinxcosx=1/2(sin2x) 2sinxcosx=sin2x

OpenStudy (anonymous):

if you need help with Trig identities go to this site it tell all the steps to get answers http://symbolab.com/solver/trigonometric-identity-calculator